OpenStudy (anonymous):

Find a number whose product with 6 is the same as its sum with 60. (Hint: you may want to try guess and check or a table.) Show all steps or reasoning. Be sure to check your answer for reasonableness.

OpenStudy (ajprincess):

Let the number be x. Then 6x=x+60 Solve for x to find the number

OpenStudy (anonymous):

doesn't work because \(10\times 6=60\) but \(60+10=70\) try 12

OpenStudy (anonymous):

THE NUMBER IS 12! haha sorry i ment to put it in the geometry sub section and when did anyone say it was 10?